Marves wants to have a total of N$4500 in two years so he can take a trip to Harare. He finds an 

account that pays 4% interest compounded monthly.

How much should Marves put into this account so that he’ll have N$4500 at the end of two 

years?


1
Expert's answer
2022-04-27T16:38:28-0400

The formula for compound interest is


"A=P(1+\dfrac{r}{n})^{nt}"

"P=A(1+\dfrac{r}{n})^{-nt}"

"P=\$4500(1+\dfrac{0.04}{12})^{-12
(2)}"

"P=\$4155"

Marves has to put $4155 into this account.